# ChromFound: Towards A Universal Foundation Model for Single-Cell Chromatin Accessibility Data

### v2

@@ Abstract
The advent of single-cell Assay for Transposase-Accessible Chromatin using sequencing (scATAC-seq) offers an innovative perspective for deciphering regulatory mechanisms by assembling a vast repository of single-cell chromatin accessibility data. While foundation models have achieved significant success in single-cell transcriptomics, there is currently no foundation model for scATAC-seq that supports zero-shot high-quality cell identification and comprehensive multi-omics analysis simultaneously. Key challenges lie in the high dimensionality and sparsity of scATAC-seq data, as well as the lack of a standardized schema for representing open chromatin regions (OCRs). Here, we present ChromFound, a foundation model tailored for scATAC-seq. ChromFound utilizes a hybrid architecture and genome-aware tokenization to effectively capture genome-wide long contexts and regulatory signals from dynamic chromatin landscapes. Pretrained on 1.97 million cells from 30 tissues and 6 disease conditions, ChromFound demonstrates broad applicability across 6 diverse tasks. Notably, it achieves robust zero-shot performance in generating universal cell representations and exhibits excellent transferability in cell type annotation and cross-omics prediction. By uncovering enhancer-gene links undetected by existing computational methods, ChromFound offers a promising framework for understanding disease risk variants in the noncoding genome.
@@ 1 Introduction
Pretrained language models (PLMs) have achieved great success in deciphering the complex language of single-cell RNA sequencing (scRNA-seq) data [67, 10, 28, 81], validating their capabilities fine-tuning for a variety of downstream tasks. Nevertheless, the development of generalizable foundation models for scATAC-seq remains under-explored, posing significant challenges for cellular representation learning and multi-task transfer. Here, we identify and discuss three major challenges.
(1). scATAC-seq data are inherently high-dimensional and sparse, with single cells spanning over a million OCRs but typically less than 1% showing accessibility. This sparsity complicates the utilization of Transformer [70] adopted in transcriptomics [67, 10, 28, 73, 19, 81]. Other methods such as transforming accessibility into gene activity [23, 62], filtering highly variable OCRs [23, 65, 21], or binarizing accessibility [29, 12, 85] always cause substantial information loss [63, 46, 5].
@@
To validate the effectiveness of ChromFound, we conduct comprehensive evaluations across multiple tasks and datasets. In cell representation, ChromFound outperforms baselines instead of additional training, achieving an average ARI improvement of 17.47% across 8 scATAC-seq datasets from 4 tissues. Additional results demonstrate its robustness to denoise technical noise in zero-shot settings. For cell type annotation, ChromFound significantly improves accuracy and macro F1 score across all evaluated datasets and tissues. In cross-omics prediction, it accurately infers gene expression from chromatin accessibility, surpassing all baselines across 5 datasets. Furthermore, in the K562 cell line, ChromFound effectively identifies gene-enhancer relationships and regulatory perturbation effects, highlighting its potential for interpretable cis-regulatory modeling.
@@ 2 Related Work
Single-cell foundation models for scRNA-seq data have seen significant advancements, supporting tasks such as gene function prediction, cell annotation, and drug response modeling. Geneformer [67], scGPT [10], and scFoundation [28] are trained by large-scale datasets, while scBERT [81] and CellPLM [73] focus on specific objectives based on pre-training language models such as imputation and cell communications. However, existing models are predominantly tailored for scRNA-seq data and fail to model scATAC-seq data effectively. GET [19] achieves a level of predictive precision comparable to experimental replicates by integrating chromatin accessibility data and genomic sequence information in a pseudo-bulk format instead of single-cell level.
Deep learning methods are widely used in scATAC-seq data analysis for cell clustering and cell annotation. SCALE [76] and its extension SCALEX [75] leverage the variational autoencoder [33] for latent feature extraction and multi-omics integration, while CASTLE [11] enhances cell type annotation through self-supervised learning. scBasset [85] connects the genomic sequence with binary accessibility and shows its strong ability to cell representation and batch integration. Cellcano  [45] and SANGO [86] employ a two-step MLP for cell type annotation. These models are impeded by their fundamental architectural limitations and the scarcity of available data, thereby restricting their scalability and generalizability across diverse datasets.
Figure 1: An Overview of ChromFound architecture. The methods of ChromFound are listed in Section 3. OCR tokenization are described in Section 3.1. The hybrid encoder layer is described in Section 3.2. Other pretraining details of datasets and implementation are included in Section 3.3. All evaluation results are detailed in Section 4.
@@ 3.1 OCR Tokenization
To encode chromatin accessibility data, we treat each OCR as a token by three essential components: chromosome embedding, positional embedding for genomic coordinates, and continuous accessibility embedding. The tokenization integrates these components to form a comprehensive representation of each OCR. In the following, we detail the encoding strategy for each component.
@@
where $p^{s}_{j}$ and $p^{e}_{j}$ represents the starting and end position for the OCR $j$ on the GRCh38 reference genome, temp is a hyperparameter set to 100000 in ChromFound, and $d_{\text{model}}$ is the dimension of the embedding layer. The index $k$ spans $\{0,1,\dots,d_{\text{model}}/2\}$, with sine and cosine functions alternately assigned to even and odd dimensions of the embedding vector.
Accessibility Embedding: The accessibility value for the $j$th OCR, denoted as $X_{j}$, is a continuous scalar processed by the procedures detailed in Appendix C. To project the value $X_{j}$ into the hidden dimension, we use a linear transformation as $\mathbf{E}_{a,j}=\text{Linear}(X_{j})$.
@@ 3.2.1 Window Partition Self-Attention (WPSA)
Due to its computational complexity $O(L^{2})$, the vanilla attention mechanism [70] is not directly applicable to scATAC-seq data containing millions of OCRs. To address these challenges, ChromFound introduces an optimized self-attention layer named Window Partition Self-Attention (WPSA), applying attention mechanism to within a window divided by window size $W\ll L$.
@@
Furthermore, we remove any zero padding in $\hat{\mathbf{E}}_{\text{wpsa}}$ and get the WPSA output $\mathbf{E}_{\text{wpsa}}\in\mathbb{R}^{L\times D}$.
@@ 3.2.2 Mamba Block
Mamba [25] is capable of processing extended sequences due to the linear-time complexity of state space models (SSMs) [27, 26], making it particularly well suited for modeling high-resolution scATAC-seq data. When combined with the WPSA module, which amplifies local signals within predefined windows, the Mamba block enables effective integration of both short- and long-range dependencies among OCRs across the genome. This hybrid architecture mitigates the degradation of long-range information typically caused by relying solely on zero-order discretization [51, 84]. To reduce parameter count and refine representations, we project $\mathbf{E}_{\text{wpsa}}$ to a lower-dimensional space $\mathbb{R}^{L\times D_{\mathrm{low}}}$ and then back, which is define as:
@@ 3.3 Pre-training
Training objective: ChromFound adopts a masking strategy that simultaneously targets both zero and non-zero OCR values. Inspired by xTrimoGene [22], the setting encourages the model to predict non-binary chromatin accessibility across the entire profile, mitigating the risk of learning representations dominated by zero entries. The MLP layer is employed as the decoder to reconstruct the original input $X_{i}$. The reconstruction output of cell $i$ is denoted as $\hat{X}_{i}$. The pretraining objective is the mean squared error (MSE) computed over the masked positions $\mathcal{M}_{i}$ in each cell $i$:
@@
Morabito130K[49] (batch 1) Cortex SCALE[76] 0.5460±0.0074 0.6896±0.0042 0.7282±0.0016 0.7275±0.0016
@@
Morabito130K[49] (batch 2) Cortex SCALE[76] 0.5306±0.0048 0.6394±0.0032 0.6221±0.0007 0.6200±0.0007
@@
Gains(%) 16.83 11.40 9.91 9.95
Kuppe139K[36] (donar av3) Heart SCALE[76] 0.4475±0.0015 0.5654±0.0011 0.6005±0.0003 0.5978±0.0003
@@
Gains(%) 28.01 17.94 10.21 10.31
Kuppe139K[36] (donar av10) Heart SCALE[76] 0.4517±0.0087 0.6596±0.0040 0.5397±0.0041 0.5391±0.0041
@@
Gains(%) 16.34 8.71 8.68 8.70
Liang154K[38] (sample D026_13) Retina SCALE[76] 0.5512±0.0386 0.7285±0.0075 0.7172±0.0051 0.7168±0.0051
@@
Gains(%) 20.98 11.86 6.59 6.60
Liang154K[38] (sample D19D008) Retina SCALE[76] 0.6129±0.0131 0.7465±0.0113 0.7981±0.0020 0.7993±0.0058
@@
PBMC169K[15] (batch VIB_10xv1_1) PBMC SCALE[76] 0.6158±0.0048 0.6959±0.0031 0.7632±0.0010 0.7623±0.0010
@@
Gains(%) 11.44 8.29 2.61 2.62
PBMC169K[15] (batch BIO_ddseq_1) PBMC SCALE[76] 0.4359±0.0181 0.6250±0.0101 0.5594±0.0045 0.5587±0.0045
@@ 4 Experiments
To systematically evaluate the performance of ChromFound, we conduct experiments on various tissues and datasets. We first assess its zero-shot representation capabilities, demonstrating strong robustness to both data sparsity and batch effects. We then evaluate the transferability in downstream tasks, including cell type annotation and cross-omics prediction. Finally, we validate the biological utility of ChromFound by accurately inferring gene-enhancer regulatory links and gene expression responses to enhancer perturbations. All scATAC-seq datasets used in experiments are preprocessed according to the procedures detailed in Appendix C.
@@ 4.1 Cell representation
ChromFound generates low-dimensional cell representations by applying PCA to the encoder outputs after the pooling layer. Details of the implementations are provided in Appendix A.1. For comparison, we include scBasset [85] and three VAE-based models [76, 75, 11], which require self-supervised learning on each specific dataset. We conduct experiments on eight datasets from four tissue types and evaluate performance using four clustering metrics to ensure fair and reliable comparison.
All results presented in Table 1 reveal that ChromFound outperforms existing SOTA methods in all tissues and metrics, with average improvements of 17.47%, 10.39%, 6.72% and 6.69% in ARI, FMI, NMI and AMI, respectively. The greater improvements observed in ARI compared to NMI and AMI infer the robustness of ChromFound to technical noise such as low read depth and batch effects, as ARI is particularly sensitive to local structures within the cell representation space. We discuss ChromFound’s ability to denoising low count and batch effect as below.
Denoising low count: Some studies [5, 37, 65] have acknowledged that the prevalence of missing signals in scATAC-seq makes standard analyses challenging. Although all datasets in Table 1 already exhibit high sparsity (~99%), we further simulate increasingly sparse conditions by downsampling accessible OCRs to 10%, 20%, 30%, 40% and 50% of their original counts. We adopt the same benchmarking methods and clustering metrics as listed in Table 1. As shown in Fig. 2, ChromFound consistently maintains stable performance across all levels of downsampling, with its advantage over baseline methods becoming increasingly pronounced as the sparsity intensifies.
Denoising batch effect: Denoising batch effects remains a major challenge in scATAC-seq analysis [47, 65, 44]. We evaluate ChromFound on datasets from 4 tissues using biological conservation metrics [44] consistent with scGPT [10]. In addition to SCALEX [75] and scBasset [85], we include five specialized batch integration methods [41, 31, 35, 77, 40] for comparison. As shown in Table 2, ChromFound in zero-shot settings outperforms existing methods by more effectively mitigating batch effects while better preserving biologically significant variation.

### v3

@@ Abstract
The advent of single-cell Assay for Transposase-Accessible Chromatin using sequencing (scATAC-seq) offers an innovative perspective for deciphering regulatory mechanisms by assembling a vast repository of single-cell chromatin accessibility data. While foundation models have achieved significant success in single-cell transcriptomics, there is currently no foundation model for scATAC-seq that supports zero-shot high-quality cell identification and comprehensive multi-omics analysis simultaneously. Key challenges lie in the high dimensionality and sparsity of scATAC-seq data, as well as the lack of a standardized schema for representing open chromatin regions (OCRs). Here, we present ChromFound, a foundation model tailored for scATAC-seq. ChromFound utilizes a hybrid architecture and genome-aware tokenization to effectively capture genome-wide long contexts and regulatory signals from dynamic chromatin landscapes. Pretrained on 1.97 million cells from 30 tissues and 6 disease conditions, ChromFound demonstrates broad applicability across 6 diverse tasks. Notably, it achieves robust zero-shot performance in generating universal cell representations and exhibits excellent transferability in cell type annotation and cross-omics prediction. By uncovering enhancer-gene links undetected by existing computational methods, ChromFound offers a promising framework for understanding disease risk variants in the noncoding genome. The implementation of ChromFound is available via https://github.com/JohnsonKlose/ChromFound.
@@ 1 Introduction
Pretrained language models (PLMs) have achieved great success in deciphering the complex language of single-cell RNA sequencing (scRNA-seq) data [64, 11, 27, 77], validating their capabilities of fine-tuning for a variety of downstream tasks. Nevertheless, the development of generalizable foundation models for scATAC-seq remains under-explored, posing significant challenges for cellular representation learning and multi-task transfer. Here, we identify and discuss three major challenges.
(1). scATAC-seq data are inherently high-dimensional and sparse, with single cells spanning over millions of OCRs but typically less than 1% showing accessibility. This sparsity complicates the utilization of Transformer [67] adopted in transcriptomics [64, 11, 27, 70, 18, 77]. Other methods such as transforming accessibility into gene activity [22, 59], filtering highly variable OCRs [22, 62, 20], or binarizing accessibility [28, 13, 81] always cause substantial information loss [60, 43, 6].
@@
To validate the effectiveness of ChromFound, we conduct comprehensive evaluations across multiple tasks and datasets. In cell representation, ChromFound outperforms baselines instead of additional training, achieving an average ARI improvement of 17.02% across 8 scATAC-seq datasets from 4 tissues. Additional results demonstrate its robustness to denoise technical effect in zero-shot settings. For cell type annotation, ChromFound significantly improves accuracy and macro F1 score across all evaluated datasets and tissues. In cross-omics prediction, it accurately infers gene expression from chromatin accessibility, surpassing all baselines across 5 datasets. Furthermore, in the K562 cell line, ChromFound effectively identifies gene-enhancer relationships and regulatory perturbation effects, highlighting its potential for interpretable cis-regulatory modeling.
@@ 2 Related Work
Single-cell foundation models for scRNA-seq data have seen significant advancements, supporting tasks such as gene function prediction, cell annotation, and drug response modeling. Geneformer [64], scGPT [11], and scFoundation [27] are trained by large-scale datasets, while scBERT [77] and CellPLM [70] focus on specific objectives based on pre-training language models such as imputation and cell communications. However, existing models are predominantly tailored for scRNA-seq data and fail to model scATAC-seq data effectively. GET [18] achieves a level of predictive precision comparable to experimental replicates. It does so by integrating chromatin accessibility data and genomic sequence information in a pseudo-bulk format, rather than relying on single-cell resolution.
Deep learning has been widely applied to scATAC-seq data analysis for cell clustering and annotation. PeakVI [1] and PoissonVI [44] extend VAE-based frameworks, with the former disentangling biological from technical variation for batch correction and the latter introducing a Poisson likelihood to preserve quantitative accessibility. SCALE [73] and SCALEX [72] employ VAEs for latent representation learning and multi-omics integration, while CASTLE [12] enhances cell-type annotation via self-supervised learning. Beyond VAE-based methods, SnapATAC2 [84] scales to millions of cells using latent semantic indexing and spectral embedding, and Signac [59] provides a Seurat-based workflow for dimensionality reduction and multimodal integration. scBasset [81] links genomic sequences with accessibility profiles, and Cellcano [42] together with SANGO [82] employ two-step MLPs for cell-type annotation. Despite their advances, these models remain limited by architectural constraints and data scarcity, restricting scalability and generalization.
Figure 1: An overview of ChromFound architecture. The methods of ChromFound are listed in Section 3. OCR tokenization are described in Section 3.1. The hybrid encoder layer is described in Section 3.2. Other pretraining details of datasets and implementation are included in Section $3.$ All evaluation results are detailed in Section 4.
@@ 3.1 OCR Tokenization
To encode chromatin accessibility data, we treat each OCR as a token by three essential components: chromosome embedding, positional embedding for genomic coordinates, and continuous accessibility embedding. The tokenization integrates these components to form a comprehensive representation of each OCR. Our unified token representation is designed to address two fundamental challenges. First, by encoding start and end genomic coordinates through sinusoidal positional embeddings, ChromFound captures the full extent of each OCR, accommodating varying fragment lengths. Second, by avoiding reliance on a fixed OCR vocabulary, ChromFound flexibly represents tissue-specific or novel OCRs, thereby mitigating misalignment and dropout across tissues. In the following, we describe the encoding strategy for each component in detail.
@@
where $p^{s}_{j}$ and $p^{e}_{j}$ represent the starting and end position for the OCR $j$ on the GRCh38 reference genome, temp is a hyperparameter set to 100000 in ChromFound, and $d_{\text{model}}$ is the dimension of the embedding layer. The index $k$ spans $\{0,1,\dots,d_{\text{model}}/2\}$, with sine and cosine functions alternately assigned to even and odd dimensions of the embedding vector.
Accessibility Embedding: The accessibility value for the $j$th OCR, denoted as $X_{j}$, is a continuous scalar processed by the procedures detailed in Appendix D. To project the value $X_{j}$ into the hidden dimension, we use a linear transformation as $\mathbf{E}_{a,j}=\text{Linear}(X_{j})$.
@@ 3.2.1 Window Partition Self-Attention (WPSA)
Due to its computational complexity $O(L^{2})$, the vanilla attention mechanism [67] is not directly applicable to scATAC-seq data containing millions of OCRs. To address these challenges, ChromFound introduces an optimized self-attention layer named Window Partition Self-Attention (WPSA), which applies the attention mechanism within local windows of size $W$, where $W\ll L$.
@@
Furthermore, we remove all zero padding tokens in $\hat{\mathbf{E}}_{\text{wpsa}}$ and get the WPSA output $\mathbf{E}_{\text{wpsa}}\in\mathbb{R}^{L\times D}$.
@@ 3.2.2 Mamba Block
Mamba [24] is capable of processing extended sequences due to the linear-time complexity of state space models (SSMs) [26, 25], making it particularly well suited for modeling high-resolution scATAC-seq data. When combined with the WPSA module, which amplifies local signals within predefined windows, the Mamba block enables effective integration of both short- and long-range dependencies among OCRs across the genome. This hybrid architecture mitigates the degradation of long-range information typically caused by relying solely on zero-order discretization [50, 80]. To reduce parameter count and refine representations, we project $\mathbf{E}_{\text{wpsa}}$ to a lower-dimensional space $\mathbb{R}^{L\times D_{\mathrm{low}}}$ and then back, which is defined as:
@@ 3.3 Pre-training
Training objective: Due to the high sparsity of scATAC-seq data, ChromFound adopts a masking strategy that simultaneously targets both zero and non-zero OCR values. Inspired by xTrimoGene [21], the setting encourages the model to predict non-binary chromatin accessibility across the entire profile, mitigating the risk of learning representations dominated by zero entries. Poisson/ZINB losses often cause unstable optimization due to gradients dominated by nonzero entries. Instead, mean squared error (MSE) loss on log-transformed, normalized signals ensures stability. The MLP layer is employed as the decoder to reconstruct the original input $X_{i}$. The reconstruction output of cell $i$ is denoted as $\hat{X}_{i}$. The pretraining objective is the MSE loss computed over the masked positions $\mathcal{M}_{i}$ in each cell $i$:
@@
Morabito130K [48] (batch 1) Cortex PeakVI [1] 0.4558±0.0038 0.6286±0.0022 0.6964±0.0012 0.6940±0.0012
PoissonVI [44] 0.3673±0.0028 0.5673±0.0018 0.5878±0.0001 0.5845±0.0001
SnapATAC2 [84] 0.5118±0.0315 0.6740±0.0145 0.7202±0.0041 0.7179±0.0042
Signac [59] 0.5013±0.0507 0.7035±0.0177 0.5580±0.0160 0.5534±0.0165
SCALE [73] 0.5460±0.0074 0.6896±0.0042 0.7282±0.0016 0.7275±0.0016
@@
Morabito130K [48] (batch 2) Cortex PeakVI [1] 0.5512±0.0016 0.6394±0.0032 0.7050±0.0006 0.7028±0.0006
PoissonVI [44] 0.5334±0.0052 0.6388±0.0033 0.6004±0.0001 0.5983±0.0001
SnapATAC2 [84] 0.5325±0.0072 0.6380±0.0045 0.7038±0.0018 0.7015±0.0018
Signac [59] 0.5069±0.0123 0.6478±0.0064 0.5888±0.0032 0.5851±0.0033
SCALE [73] 0.5306±0.0048 0.6394±0.0032 0.6221±0.0007 0.6200±0.0007
@@
Gains(%) 16.83 10.47 3.84 3.86
Kuppe139K [33] (donar av3) Heart PeakVI [1] 0.4159±0.0007 0.5434±0.0006 0.6024±0.0005 0.6000±0.0006
PoissonVI [44] 0.3969±0.0005 0.5232±0.0004 0.5717±0.0001 0.5688±0.0001
SnapATAC2 [84] 0.4571±0.0099 0.5564±0.0064 0.6625±0.0013 0.6502±0.0013
Signac [59] 0.3123±0.0047 0.4601±0.0028 0.5272±0.0028 0.5238±0.0029
SCALE [73] 0.4475±0.0015 0.5654±0.0011 0.6005±0.0003 0.5978±0.0003
@@
Gains(%) 27.75 17.94 8.78 10.31
Kuppe139K [33] (donar av10) Heart PeakVI [1] 0.4301±0.0030 0.6383±0.0021 0.6149±0.0022 0.6144±0.0022
PoissonVI [44] 0.3606±0.0054 0.5929±0.0027 0.4472±0.0066 0.4464±0.0066
SnapATAC2 [84] 0.5515±0.0312 0.7285±0.0124 0.6867±0.0060 0.6863±0.0060
Signac [59] 0.1159±0.0112 0.5093±0.0053 0.2294±0.0076 0.2281±0.0076
SCALE [73] 0.4517±0.0087 0.6596±0.0040 0.5397±0.0041 0.5391±0.0041
@@
Gains(%) 16.34 8.71 7.31 7.31
Liang154K [35] (sample D026_13) Retina PeakVI [1] 0.2885±0.0024 0.5543±0.0019 0.5644±0.0019 0.5638±0.0019
PoissonVI [44] 0.3162±0.0025 0.5766±0.0017 0.6081±0.0017 0.6076±0.0017
SnapATAC2 [84] 0.5606±0.0569 0.7207±0.0199 0.7045±0.0085 0.7041±0.0085
Signac [59] 0.5538±0.1051 0.7451±0.0095 0.4931±0.0535 0.4919±0.0538
SCALE [73] 0.5512±0.0386 0.7285±0.0075 0.7172±0.0051 0.7168±0.0051
@@
Gains(%) 18.94 9.37 6.59 6.60
Liang154K [35] (sample D19D008) Retina PeakVI [1] 0.4702±0.0061 0.6318±0.0037 0.7229±0.0024 0.7224±0.0025
PoissonVI [44] 0.4978±0.0122 0.6527±0.0065 0.7400±0.0028 0.7395±0.0028
SnapATAC2 [84] 0.5814±0.0019 0.7173±0.0010 0.7881±0.0011 0.7877±0.0011
Signac [59] 0.5146±0.0532 0.7339±0.0119 0.5235±0.0307 0.5222±0.0309
SCALE [73] 0.6129±0.0131 0.7465±0.0113 0.7981±0.0020 0.7993±0.0058
@@
PBMC169K [15] (batch VIB_10xv1_1) PBMC PeakVI [1] 0.5417±0.0022 0.6340±0.0015 0.6639±0.0009 0.6627±0.0009
PoissonVI [44] 0.6314±0.0027 0.7073±0.0018 0.7353±0.0012 0.7343±0.0012
SnapATAC2 [84] 0.6246±0.0112 0.7085±0.0067 0.7551±0.0023 0.7541±0.0023
Signac [59] 0.1832±0.0136 0.4543±0.0043 0.3064±0.0198 0.3027±0.0201
SCALE [73] 0.6158±0.0048 0.6959±0.0031 0.7632±0.0010 0.7623±0.0010
@@
Gains(%) 10.12 7.28 2.61 2.62
PBMC169K [15] (batch BIO_ddseq_1) PBMC PeakVI [1] 0.2915±0.0005 0.5041±0.0004 0.4623±0.0004 0.4615±0.0005
PoissonVI [44] 0.4103±0.0045 0.5994±0.0028 0.5656±0.0035 0.5649±0.0035
SnapATAC2 [84] 0.4126±0.0318 0.6149±0.0149 0.5761±0.0071 0.5754±0.0071
Signac [59] 0.1885±0.0097 0.5249±0.0056 0.2852±0.0040 0.2839±0.0040
SCALE [73] 0.4359±0.0181 0.6250±0.0101 0.5594±0.0045 0.5587±0.0045
@@ 4 Experiments
To systematically evaluate the performance of ChromFound, we conduct experiments on various tissues and datasets. We first assess its zero-shot representation capabilities, demonstrating strong robustness to both data sparsity and batch effects. We then evaluate the transferability in downstream tasks, including cell type annotation and cross-omics prediction. Finally, we validate the biological utility of ChromFound by accurately inferring gene-enhancer regulatory links and gene expression responses to enhancer perturbations. All scATAC-seq datasets used in experiments are preprocessed according to the procedures detailed in Appendix D.
@@ 4.1 Cell representation
ChromFound generates low-dimensional cell representations by applying PCA to the encoder outputs after the pooling layer. Details of the implementations are provided in Appendix B.1. For comparison, we include scBasset [81] and three VAE-based models [73, 72, 12], which require self-supervised learning on each specific dataset. We conduct experiments on eight datasets from four tissue types and evaluate performance using four clustering metrics to ensure a fair and reliable comparison.
All results presented in Table 1 reveal that ChromFound outperforms existing SOTA methods in all tissues and metrics, with average improvements of 17.02%, 10.39%, 6.72% and 6.69% in ARI, FMI, NMI and AMI, respectively. The greater improvements observed in ARI compared to NMI and AMI infer the robustness of ChromFound to technical noise such as low read depth and batch effects, as ARI is particularly sensitive to local structures within the cell representation space. We discuss ChromFound’s ability to denoise low count and batch effect as below.
Denoising low count: Some studies [6, 34, 62] have acknowledged that the prevalence of missing signals in scATAC-seq makes standard analyses challenging. Although all datasets in Table 1 already exhibit high sparsity (~99%), we further simulate increasingly sparse conditions by downsampling accessible OCRs to 10%, 20%, 30%, 40% and 50% of their original counts. We adopt the same benchmarking methods and clustering metrics as listed in Table 1. As shown in Fig. 2, ChromFound consistently maintains stable performance across all levels of downsampling, with its advantage over baseline methods becoming increasingly pronounced as the sparsity intensifies. Furthermore, we observe that ChromFound performs better on the retina dataset under stronger downsampling. This improvement likely reflects protocol-specific artifacts, as retina data contain more frequently accessible OCRs, introducing redundancy and noise. Random downsampling suppresses these excessive signals, enhances the signal-to-noise ratio, and yields the observed performance gain.
@@
Denoising batch effect: Denoising batch effects remains a major challenge in scATAC-seq analysis [46, 62, 41]. We evaluate ChromFound on datasets from 4 tissues using biological conservation metrics [41] consistent with scGPT [11]. In addition to PeakVI [1], PoissonVI [44], SCALEX [72] and scBasset [81], we include five specialized batch integration methods [38, 29, 32, 74, 37] for comparison. As shown in Table 2, ChromFound in zero-shot settings outperforms existing methods by more effectively mitigating batch effects while better preserving biologically significant variation.